OpenStudy (lolipop_drum312):

How many minutes equal 5.4 days? Enter your answer in the box. 5.4 days = min (7776 is my answer am i right?)

OpenStudy (mathmale):

Please show your work if you want feedback on it. Also, your answer should have a unit of measurement; 7776 can't stand alone and still make sense here.

OpenStudy (lolipop_drum312):

well i was thinking what if i divided it by something

OpenStudy (lolipop_drum312):

idk

OpenStudy (rcp031):

U must have an understanding of unit of conversion before solving such question as mentioned by @mathmale

OpenStudy (lolipop_drum312):

oh ok...... i just wanted to see if i'm right

OpenStudy (lolipop_drum312):

@Emeri_Renteria

OpenStudy (emeri_renteria):

im seeing

OpenStudy (lolipop_drum312):

ok XD

OpenStudy (lolipop_drum312):

so any ideas

OpenStudy (emeri_renteria):

ok so first you have to see how many min are in a days

OpenStudy (emeri_renteria):

lol

OpenStudy (lolipop_drum312):

but its the samee 5.4 right unless its differ

OpenStudy (emeri_renteria):

hmmmmmm.....

OpenStudy (rcp031):

unit of conversion 1day= 1440 minutes or 1 minute =1/1440 day

OpenStudy (lolipop_drum312):

i mean 776 up top

OpenStudy (lolipop_drum312):

oh so that will still be um 1440 a day if we divided those 2 right

OpenStudy (lolipop_drum312):

i mean 1

OpenStudy (emeri_renteria):

ok now you can multiply 1440 times 5 days

OpenStudy (lolipop_drum312):

ok and we get 7200 by multiplying those 2 right

OpenStudy (emeri_renteria):

yes

OpenStudy (emeri_renteria):

wait ..?

OpenStudy (lolipop_drum312):

hmm? 0.0

OpenStudy (emeri_renteria):

im seeing

OpenStudy (lolipop_drum312):

ok

OpenStudy (emeri_renteria):

yes thats in 5 days

OpenStudy (emeri_renteria):

7200 mins

OpenStudy (lolipop_drum312):

so our answer is 7200

OpenStudy (emeri_renteria):

no but wait you gotta do the .4

OpenStudy (emeri_renteria):

so you we said that 7200 mins in five days, ok then you have to divide now

OpenStudy (emeri_renteria):

since theres 1440 in one day divide it by four

OpenStudy (lolipop_drum312):

360

OpenStudy (mathmale):

You need to learn some basics. First, 60 minutes = 1 hour; 24 hours = 1 day. From this info you can obtain 1440 minutes = 1 day. This info was given to you by another person in this and in your previous problem. I'd suggest you WRITE DOWN info that you want to remember. It would have saved you so much time and trouble here.

OpenStudy (lolipop_drum312):

i do write it down

OpenStudy (lolipop_drum312):

but i have to understand it first

OpenStudy (lolipop_drum312):

then i work it out

OpenStudy (mathmale):

I'd be happy to give you feedback on your work if you'd share what you've done. An "is that right?" question doesn't bring in ymuch info. Fine. Show how you start with 60 min = 1 hour and 24 hours = 1 day and come up with 1440 hours = 1 day. Can you do that?

OpenStudy (lolipop_drum312):

but i don't know how to do all that thats why i'm trying to work it out with all of you so i can get it....

OpenStudy (lolipop_drum312):

do i divide all of it.. what do i do help me work it....

OpenStudy (lolipop_drum312):

@Twistrose

OpenStudy (twistrose):

ok so what do you have so far

OpenStudy (lolipop_drum312):

well we divided and got 360

OpenStudy (lolipop_drum312):

you got an idea

OpenStudy (mathmale):

First we'll change "days" into "hours." The relevant equality is 24 hours = 1 day. Then,\[\frac{ 5.4~days }{ 1 }\frac{ 24~hours }{ 1~day}=?\]